OpenStudy (faman39):

Peter earned a wage increase of 15% . If he was previously making 9% per hour, how much more is Peter making per hour now?

OpenStudy (amistre64):

9% per hour? or $9 per hour?

OpenStudy (amistre64):

cause otherwise it reads as tho hes making 15% per hour now .....

OpenStudy (saifoo.khan):

it must be $9 @Faman39

OpenStudy (saifoo.khan):

So, we will find 15% of $9 and add it to $9 to get the new wage.

OpenStudy (saifoo.khan):

\[9 * \frac{15}{100}\]\[$1.35\]

OpenStudy (saifoo.khan):

Therefore, your total price = 9 + 1.35 = $10.35

OpenStudy (saifoo.khan):

Your name's fatima! D: You are a Muslim? :D